# Continual Action Assessment via Task-Consistent Score-Discriminative Feature Distribution Modeling

### v4

@@ IV-C Action General-Specific Graph
Considering the shared and specific criteria across various AQA tasks, we build our general-specific knowledge Modeling module, namely Action General-Specific Graph (AGSG), on JRG [4] to learn and decouple action-general and action-specific knowledge so that task-consistent features can be better extracted. The pipeline of AGSG can be shown in Figure 4.
@@
$A^{t}=\alpha*G+(1-\alpha)*S^{t},$ (7)
where $\alpha\in[0,1]$ is a hyperparameter, $G$ denotes the AGG and $S^{t}$ denotes the ASG.
@@
Here we introduce the subscript $g\in\{1,...,t-1\}$ for $E^{t}$ and $E^{t-1}$ to denote that $A^{g}$ is utilized to extract features. Although those features extracted when using AGSGs except $A^{\epsilon}$ are meaningless for the inference stage, inspired by LwF [37], we believe they are valuable as pseudo labels for distillation.
Noting that it is essential to keep the model initialized in an identical way for fair comparisons, and a better initialization can help a model learn better. Hence, at the base step ($t\text{=}1$), we initialize the ASG to keep the initial $A^{1}$ the same as those in the model without ASG. At continual training steps, we initialize $S^{t}$ by the previously learnt ASG at task $t\text{-}1$. Experiments in Sec. V-E show that such initialization can gain performance improvement.
- Discussions. We develop a stronger feature extractor rather than directly adopting the original JRG [4] for the following two reasons. First, the importance of the relationship between joint pairs may differ among actions, for which using a single graph to model joint relationships across different actions is not feasible for Continual-AQA. Second, considering the general and specific judging criteria across different tasks, action-general and action-specific knowledge should be explicitly modeled. Hence, AGSG is proposed to model such dual knowledge during continual training. We believe modeling and decoupling the general and specific knowledge across various tasks would further mitigate forgetting because the general knowledge can be shared, and the specific knowledge can be isolated so that task-consistent features can be better extracted. Notably, the AGSG module has two-fold novelties: 1) Compared with the existing works on AQA, AGSG is the first module explicitly models general and specific knowledge. 2) Compared with the original JRG, AGSG is a CL-based module that differs from JRG in graph calculation and model training.
@@ IV-D Continual Training Strategy
As introduced in IV-B, FS-Aug directly augments $f_{pre,j}^{t-1}$ and $s_{pre,j}$. Therefore, instead of mixing the previous data and current data together, given a mini-batch size $b$ for current data, we additionally sample a mini-batch from previous data to compute $f_{pre,j}^{t}$ and $f_{pre,j}^{t-1}$. Moreover, during training, given an instance $v^{\epsilon}_{i}$ from task $\epsilon$, $f^{t}$ and $f^{t-1}$ introduced in Section IV-B are extracted using the AGSG $A^{\epsilon}$.
@@
where $\lambda_{FD}$ and $\lambda_{Diff}$ are hyper-parameters. Following previous works [4, 26], AQA loss is defined as the MSE loss across the predicted and ground-truth scores. Note that when training the model on the first task, the later two losses are not computed to optimize the model. For more training details, please refer to Algorithm 1.
@@ V-A Datasets and Evaluation Metrics
MTL-AQA [26]: The MTL-AQA dataset contains 1,412 video actions from 16 different world diving events. As for the annotations, the MTL-AQA contains various annotations, including difficulty, action type, scores from each judge, final scores, and video captions. Only the final scores are used in our work.
@@
- Common AQA Metrics. In previous works [35, 21, 6, 12, 11], Pairwise Accuracy and Spearman’s rank correlation coefficient (SRCC) are utilized to evaluate the performance of the model on an AQA task. The Pairwise Accuracy represents the percentage of correctly ranked pairs. The SRCC can evaluate the correlation between the scores series predicted by the model and the ground-truth scores series, which is defined as follows:
@@ V-B Implementation Details
For the MTL-AQA dataset, we separate it into two subsets (i.e., 3m-springboard and 10m-platform) based on the type of action, which means the number of tasks is 2. Note that under this setting, in the experiments, the values of $\mathcal{NBT}$ are always equal to the values of $\mathcal{MF}$. Hence, in the main paper, only the results on $\mathcal{AP}$ and $\mathcal{NBT}$ are reported in Table 8. We follow [31] to extract I3D [1] features from raw videos and optical flows in advance.
@@
- Model Details. For experiments on the AQA-7 dataset, We follow [4] to build our model except for the graph calculation. We separate the original model into two parts: a feature extractor and a score regressor. The difference regressor in our model has a similar architecture as the score regressor except for the input dimension.
@@
- Training Settings. In our work, Adam [40] is used to optimize our model. We set the learning rate as 0.01 for AGG and ASG and 0.001 for the other parameters in our model with a $10^{-5}$ weight decay. For our Grouping Sampling, we sample the first exemplar for the former $m\text{-}1$ groups and the last exemplar for the $m$-th group to cover the highest and lowest scores. We set the total memory size $M$ to 30 unless specified otherwise. Following iCaRL [16], we gradually remove exemplars in memory to store new exemplars. To mitigate the influence of specific task orders, for experiments on the AQA-7 and the BEST dataset, we use four random seeds [^1] Random Seeds: 0, 1, 2, 3 to generate four task sequences. For the MTL-AQA dataset, we conduct experiments on two possible task sequences. The reported results are average performance on different task sequences.
- Existing CL Baselines. In Section V-F, we compare the proposed method with four existing CL methods, including EWC [22], iCaRL [16], PODNet [24] and AFC [23]. The introduction and processing details of the baseline models are as follows:
@@
- Existing AQA Baselines. In Section V-F, we also compare the proposed method with two more existing AQA methods(i.e., USDL [25] and CoRe-GART [21]). The introduction and processing details of the baseline models are as follows:
USDL [25]: USDL tries to address AQA by reformulating it as a Distribution Prediction problem. In this work, we implement two variants: 1) original USDL, which follows the official implementation of USDL to train the model. 2) USDL*, whose I3D backbone is frozen, and only the distribution predictor is trained across tasks.
CoRe-GART [21]: CoRe-GART proposes to address AQA by reformulating it as a Contrastive Regression Problem, that is, regressing the relative scores between query videos and exemplar videos. In this work, we implement CoRe-GART by following the official implementation of it.
@@ V-D Analysis on Feature-Score Correlation-Aware Rehearsal
Figure 5: T-SNE Visualizations of the features sampled by Herding (left) and our GS (right) after training on the first task. Points in greens denotes the features. Darker color indicates higher score. Points in orange and red denote the feature sampled by Herding and our GS, respectively. The exemplars sampled by our GS can cover all score levels and roughly describe the learned feature distribution rather than gathering in a compact region. Best viewed in color.
@@
- Effectiveness of Grouping Sampling. In our work, the proposed rehearsal-based method is based on the Grouping Sampling strategy. Note that lots of methods [24, 19, 41] directly adopt the Herding strategy [16] to sample representative exemplars. We conduct a comparison among Herding, Random Sampling (RS) and our Grouping Sampling (GS) under two different memory sizes, i.e., 30 and 60. As shown in Table IV, Herding gets the worst results, even worse than Random Sampling, and our GS strategy outperforms other methods regardless of the memory size. The results suggest that Herding cannot select the most representative exemplars under Continual-AQA settings. The visualization results shown in Figure 5 also support this view. The features sampled by Herding are gathered in a compact region in the latent space. The exemplars’ corresponding scores are pretty close, which cannot honestly describe the previously learned feature distribution. On the contrary, our GS, which selects representative exemplars to roughly describe the previously learned feature distribution, can serve as a general sampling strategy for Continual-AQA.
@@
Moreover, we study the impact of the number of helpers. As shown in Figure 6, the performance improves when the number of helpers grows from 3 to 7. However, the performance is relatively poor when further increasing the number of helpers from 7 to 9. It suggests that when we set an extremely low value to the number of helpers, the augmentation direction obtained by the helpers may not be reliable enough. When an extremely high value is set to be the number of helpers, the impact of each helper tends to be offset due to the opposite direction of the perturbations.
@@
TABLE VI: Ablation studies of the initialization for Action-specific Graph and the reformulated feature distillation loss on the AQA-7 dataset.
@@ V-E Analysis on Action General-Specific Graph
In our work, an Action General-Specific Graph module is developed to learn and decouple the action-general and action-specific knowledge during continual training. As introduced in Section IV-C, we utilize specific initialization for Action-Specific Graph. Moreover, we design a specific feature distillation loss to save and decouple action-general and action-specific knowledge. To demonstrate the effectiveness of such designs, on the one hand, we replace the proposed initialization with two intuitive strategies (i.e., zero initialization and random initialization). On the other hand, we replace the proposed feature distillation loss with the naive feature distillation loss in Equation 14. As shown in Table VI, when ablating any of them, the performance declines significantly, which verifies the contributions of such designs to the performance.
@@ V-F Comparison with the Existing Methods on the AQA-7 dataset
TABLE VII: Comparisons of our approach with existing AQA methods on the AQA-7 dataset. “w/ Our GS” denotes our Grouping Sampling is adopted to sample the exemplars. “w/ PODLoss” denotes the model is trianed together with the POD Loss [24]. USDL* indicates that we frozen the I3D backbone when training USDL model.
@@
TABLE VIII: Comparisons of our approach with existing CL methods on the AQA-7 dataset. “w/ Our GS” denotes our Grouping Sampling is adopted to sample the exemplars.
Finetune 0.5523 0.2515 0.3420
Upper Bound 0.7344 - -
@@
PODNet [24] w/ Our GS 0.6393 0.1371 0.1987
AFC [23] w/ Our GS 0.6489 0.1275 0.1806
Ours 0.6492 0.1085 0.1399
TABLE IX: Comparison results with various total memory sizes (M) on the AQA-7 dataset. “w/ Our GS” indicates our Grouping Sampling is adopted to sample the exemplars.
@@
- Comparison with existing AQA methods. We build Cl baselines based on successful AQA models (i.e., USDL[25], CoRe-GART[21]) and compare them with our approach. The comparison results are shown in Table VII, and we have the following observations. First, when directly fine-tuning models across tasks, each method performs poorly with high forgetting. Second, we train USDL and CoRe-GART to build stronger baselines with the POD loss [24] and our GS. However, although POD loss and GS significantly improve the performance of USDL and CoRe-GART, our approach outperforms these two methods.
It is worthwhile noting that the enormous contrast in performance of USDL and CoRe-GART on conventional AQA tasks [25, 21] and Continual-AQA may lie in the great task-specific knowledge learning ability with the special problem formulations (e.g., Uncertain Distribution Prediction and Contrastive Regression). The more effective to learn specific task, the more easy to forget the already learned knowledge.
- Comparison with existing CL methods. We compare the proposed method with four existing CL methods (i.e, EWC[22], icarl[16], podnet[24], and AFC[23]). As shown in Table VIII, our method bridges the gap between “Finetune” and “Upper Bound” on $\mathcal{AP}$ and significantly alleviates forgetting. Additionally, our method outperforms other methods among the three metrics. It has been discussed in Section V-D that Herding may not be an appropriate sampling strategy for Continual-AQA. Hence, we further adopt our Grouping Sampling on PODNet and AFC for better comparison. As shown in Table VIII, when our Grouping Sampling is adopted, existing CL methods can perform better but still underperform our approach, further showing our approach’s effectiveness.
- Robustness study on memory size. We further evaluate the robustness of our approach by adjusting the total memory size (M). As shown in Table IX, our approach outperforms PODNet [24] regardless of the memory size, even if our GS is adopted.
Figure 7: Visualizations of the performance matrix $P$ learned by EWC [22], iCaRL [16], PODNet [24], our naive model, our GS, our FSCAR and our full model on AQA-7 dataset. The i-row-j-column element of $P$ denotes the Spearman’s rank correlation coefficient on the $i$-th task after learning the $j$-th task. Darker color indicates better performance. The task sequence is: Sync.10m, Skiing, Gym vault, Snowboard, Diving, Sync.3m.
TABLE X: Comparisons of our approach with existing Continual Learning methods on the BEST dataset. “w/ Our GS” indicates our Grouping Sampling is adopted to sample the exemplars.
@@
TABLE XI: Comparisons of our approach with existing Continual Learning methods on the MTL-AQA dataset. “w/ Our GS” indicates our Grouping Sampling is adopted to sample the exemplars.

### v5

@@ IV-C Action General-Specific Graph
As discussed in Section I, different AQA tasks may have different assessment criteria, which poses a challenge to extracting task-consistent score-discriminative features. Inspired by Complementary Learning Systems [37, 38], we believe if we can use a shared action-general module to model shared information across tasks and several specific modules for each task to model action-specific information, task-consistent score-discriminative features can be easier to extract and the forgetting in Continual-AQA will be further mitigated. Therefore, we build our general-specific knowledge Modeling module, termed the Action General-Specific Graph (AGSG), on JRG [6] to learn and decouple action-general and action-specific knowledge so that task-consistent features can be better extracted. The pipeline of AGSG can be shown in Figure 4.
@@
$A^{t}=\alpha\cdot G+(1-\alpha)\cdot S^{t},$ (7)
where $\alpha\in[0,1]$ is a hyperparameter, $G$ denotes the AGG and $S^{t}$ denotes the ASG. We will provide more details about feature extractor with AGSG in the Appendix.
@@
Here we introduce the subscript $g\in\{1,...,t-1\}$ for $E^{t}$ and $E^{t-1}$ to represent that $A^{g}$ is utilized to extract features. Although those features extracted when using AGSGs except $A^{\epsilon}$ are meaningless for the inference stage, inspired by LwF [41], we believe they are valuable as pseudo labels for distillation.
Notably, it is essential to keep the model initialized in an identical way for fair comparisons, and a better initialization can help a model learn better. Hence, at the base step ($t\text{=}1$), we initialize the ASG to keep the initial $A^{1}$ the same as those in the model without ASG. At continual training steps, we initialize $S^{t}$ by the previously learnt ASG at task $t\text{-}1$. Experiments in Sec. V-E show that such initialization can gain performance improvement.
- Discussions. We develop a stronger feature extractor rather than directly adopting the original JRG [6] for the following two reasons. First, the importance of the relationship between joint pairs may differ among actions, for which using a single graph to model joint relationships across different actions is not feasible for Continual-AQA. Second, considering the general and specific judging criteria across different tasks, action-general and action-specific knowledge should be explicitly modeled. Hence, AGSG is proposed to model such dual knowledge during continual training. We believe modeling and decoupling the general and specific knowledge across various tasks would further mitigate forgetting because the general knowledge can be shared, and the specific knowledge can be isolated so that task-consistent features can be better extracted. Notably, the AGSG module has two-fold novelties: 1) Compared with the existing works on AQA, AGSG is the first module explicitly models general and specific knowledge. 2) Compared with the original JRG, AGSG is a CL-based module that differs from JRG in graph construction and model training.
@@ IV-D Continual Training Strategy
As introduced in IV-B, FS-Aug directly augments $f_{pre,j}^{t-1}$ and $s_{pre,j}$. Therefore, instead of mixing the previous and current data together, given a mini-batch size $b$ for current data, we additionally sample a mini-batch from previous data to compute $f_{pre,j}^{t}$ and $f_{pre,j}^{t-1}$. Moreover, during training, given an instance $v^{\epsilon}_{i}$ from task $\epsilon$, $f^{t}$ and $f^{t-1}$ introduced in Section IV-B are extracted using the AGSG $A^{\epsilon}$.
@@
where $\lambda_{FD}$ and $\lambda_{Diff}$ are hyper-parameters. Following previous works [6, 28], AQA loss is defined as the MSE loss across the predicted and ground-truth scores. Note that when training the model on the first task, the later two losses are not computed to optimize the model. For more training details, please refer to the Appendix.
@@ V-A Datasets and Evaluation Metrics
  • MTL-AQA [28]: The MTL-AQA dataset contains 1,412 videos from 16 different world diving events. As for the annotations, the MTL-AQA contains various annotations, including difficulty, action type, scores from each judge, final scores, and video captions. Only the final scores are used in our work.
@@
- Common AQA Metrics. In previous works [40, 23, 8, 14, 13], Pairwise Accuracy and Spearman’s rank correlation coefficient (SRCC) are utilized to evaluate the performance of the model on an AQA task. The Pairwise Accuracy represents the percentage of correctly ranked pairs. The SRCC can evaluate the correlation between the score series predicted by the model and the ground-truth score series, which is defined as follows:
@@ V-B Implementation Details
For the MTL-AQA dataset, we separate it into two subsets (i.e., 3m-springboard and 10m-platform) based on the type of action, which means the number of tasks is 2. Note that under this setting, in the experiments, the values of $\mathcal{NBT}$ are always equal to the values of $\mathcal{MF}$. Hence, only the results on $\mathcal{AP}$ and $\mathcal{NBT}$ are reported in Table 8. We follow [34] to extract I3D [1] features from raw videos and optical flows in advance.
@@
- Model Details. For experiments on the AQA-7 dataset, We follow [6] to build our model except for the graph construction. We separate the original model into two parts: a feature extractor and a score regressor. The difference regressor in our model has a similar architecture as the score regressor except for the input dimension.
@@
- Training Settings. In our work, Adam [45] is used to optimize our model. We set the learning rate as 0.01 for AGG and ASG and 0.001 for the other parameters in our model with a $10^{-5}$ weight decay. For our Grouping Sampling, we sample the first exemplar for the former $m\text{-}1$ groups and the last exemplar for the $m$-th group to cover the highest and lowest scores. We set the total memory size $M$ to 30 unless specified otherwise. Note that such a memory size will not incur significant storage overhead because the I3D backbone is frozen during training, we can directly store the I3D features and score labels of the instances (e.g., it roughly requires a storage overhead of 41MB in our experiments on the AQA-7 datasets). Even if we store the raw videos, 30 videos will not bring significant storage overhead. Following iCaRL [18], we gradually remove exemplars in memory to store new exemplars. To mitigate the influence of specific task orders, for experiments on the AQA-7 and the BEST dataset, we use four random seeds [^1] Random Seeds: 0, 1, 2, 3 to generate four task sequences. For the MTL-AQA dataset, we conduct experiments on two possible task sequences. The reported results are average performance on different task sequences.
- Existing CL Baselines. In Section V-F, we compare the proposed method with four existing CL methods, including EWC [24], iCaRL [18], PODNet [26] and AFC [25]. The processing details of the baseline models are as follows:
@@
- Existing AQA Baselines. In Section V-F, we also compare the proposed method with two more existing AQA methods (i.e., USDL [27] and CoRe-GART [23]). The introductions and processing details of the baseline models are as follows:
  • USDL [27]: USDL tries to address AQA by reformulating it as a Distribution Prediction problem. In this work, we implement two variants: 1) original USDL, which follows the official implementation to train the model. 2) USDL*, whose I3D backbone is frozen, and only the distribution predictor is trained across tasks.
  • CoRe-GART [23]: CoRe-GART proposes to address AQA by reformulating it as a Contrastive Regression Problem, that is, regressing the relative scores between query videos and exemplar videos. In this work, we implement CoRe-GART by following the official implementation.
@@ V-D Analysis on Feature-Score Correlation-Aware Rehearsal
Figure 5: T-SNE Visualizations of the features sampled by Herding (a), Random Sampling (b) and our GS (c) after training on the first task. Points in greens denote the features. Darker color indicates higher score. Points in orange, blue and red denote the feature sampled by Herding, Random Sampling and our GS, respectively. The exemplars sampled by our GS can cover all score levels and roughly describe the learned feature distribution rather than gathering in a compact region. Best viewed in color.
@@
- Effectiveness of Grouping Sampling. In our work, the proposed rehearsal-based method is based on the Grouping Sampling strategy. Note that lots of methods [26, 21, 46] directly adopt the Herding strategy [18] to sample representative exemplars. We conduct a comparison among Herding, Random Sampling (RS) and our Grouping Sampling (GS) under two different memory sizes, i.e., 30 and 60. As shown in Table IV, when adopting different sampling strategies on our naive model (i.e., Row 1 in Table I), Herding gets the worst results, even worse than Random Sampling, and our GS strategy outperforms other methods regardless of the memory size. The results suggest that Herding cannot select the most representative exemplars under Continual-AQA settings. The visualization results shown in Figure 5 also support this view. The features sampled by Herding Figure 5(a) are gathered in a compact region in the latent space. The exemplars’ corresponding scores are pretty close, which cannot honestly describe the previously learned feature distribution. Additionally, although features with various levels of scores could be sampled by random sampling, still, the real data distribution cannot be accurately described, i.e., some exemplars with pretty close scores will be sampled. Also, data with some levels of scores may be missed by RS, i.e, in Figure 5(b) no data with high scores is sampled by RS. On the contrary, our GS Figure 5(c) can cover all score levels and roughly describe the learned feature distribution rather than gathering in a compact region, which can serve as a general sampling strategy for Continual-AQA.
@@
- Impact of the selection of helpers. We first study the impact of the number of helpers. As shown in Figure 6, the performance improves when the number of helpers grows from 3 to 7. It suggests that when we set an extremely low value to the number of helpers, the perturbation direction obtained by the helpers may not be reliable enough. In another words, it is necessary to model the feature-score relations between the feature to be augmented and the helpers with various scores. Moreover, the performance achieves slightly decline when further increasing the number of helpers from 7 to 9. We think this is because when an extremely high value is set to be the number of helpers, the impact of each helper tends to be counteracted due to the opposite direction of the perturbations.
Notably, we use the simplest solution, randomly sampling (noted as Random), to select helpers by default. We beliveve it is desirable to explore if there is a better way to select helpers. To inspire future works, we develop an Anchor-based helper selection strategy. We introduce the details of such strategy in Appendix. As shown in Table VI, the performance gains improvement with the Anchor-based helper selection strategy.
Although “Anchor-based” shows its effectiveness, it requires multiple ranking phases during training, which brings further time cost. Therefore, it is still meaningful to explore better strategies. Also, different helpers may have different importance for augmenting features and scores. We hope future effort will be devoted to these issues to further mitigate the forgetting in Continual-AQA.
TABLE VI: Comparison between different helper selection strategies.
Random 0.6492 0.1085 0.1399
Anchor-based 0.6545 0.0903 0.1277
@@
TABLE VII: Ablation studies of the initialization for Action-specific Graph and the reformulated feature distillation loss on the AQA-7 dataset.
@@ V-E Analysis on Action General-Specific Graph
In our work, an Action General-Specific Graph module is developed to learn and decouple the action-general and action-specific knowledge during continual training. As introduced in Section IV-C, we utilize specific initialization for Action-Specific Graph. Moreover, we design a specific feature distillation loss to save and decouple action-general and action-specific knowledge. To demonstrate the effectiveness of such designs, on the one hand, we replace the proposed initialization with two intuitive strategies (i.e., zero initialization and random initialization). On the other hand, we replace the proposed feature distillation loss with the naive feature distillation loss in Equation 14. As shown in Table VII, when ablating any of them, the performance declines significantly, which verifies the contributions of such designs to the performance.
@@ V-F Comparison with the Existing Methods on the AQA-7 dataset
TABLE VIII: Comparisons of our approach with existing AQA methods on the AQA-7 dataset. “w/ Our GS” denotes our Grouping Sampling is adopted to sample the exemplars. “w/ PODLoss” denotes the model is trianed together with the POD Loss [26]. USDL* indicates that we frozen the I3D backbone when training USDL model.
@@
TABLE IX: Comparisons of our approach with existing CL methods on the AQA-7 dataset. “w/ Our GS” denotes our Grouping Sampling is adopted to sample the exemplars. “# Param(m)” indicates the number of trainable parameters.
Methods # Param(m) $\mathcal{AP}(\uparrow)$ $\mathcal{NBT}(\downarrow)$ $\mathcal{MF}(\downarrow)$
Finetune 1.265 0.5523 0.2515 0.3420
Upper Bound 1.265 0.7344 - -
@@
PODNet [26] w/ Our GS 1.265 0.6393 0.1371 0.1987
AFC [25] w/ Our GS 1.265 0.6489 0.1275 0.1806
Ours 1.279 (+1.17%) 0.6492 0.1085 0.1399
TABLE X: Comparison results with various total memory sizes (M) on the AQA-7 dataset. “w/ Our GS” indicates our Grouping Sampling is adopted to sample the exemplars.
@@
- Comparison with existing AQA methods. We build CL baselines based on successful AQA models (i.e., USDL [27], CoRe-GART [23]) and compare them with our approach. The comparison results are shown in Table VIII, and we have the following observations. First, when directly fine-tuning models across tasks, each method performs poorly with high forgetting. Second, we train USDL and CoRe-GART to build stronger baselines with the POD loss [26] and our GS. However, although POD loss and GS significantly improve the performance of USDL and CoRe-GART, our approach outperforms these two methods.
It is worthwhile noting that the enormous contrast in performance of USDL and CoRe-GART on conventional AQA tasks [27, 23] and Continual-AQA may lie in the great task-specific knowledge learning ability with the special problem formulations (e.g., Uncertain Distribution Prediction and Contrastive Regression). The more effective to learn specific task, the easier to forget the already learned knowledge.
- Comparison with existing CL methods. We compare the proposed method with four existing CL methods (i.e, EWC [24], icarl [18], podnet [26], and AFC [25]). As shown in Table IX, our method bridges the gap between “Finetune” and “Upper Bound” on $\mathcal{AP}$ and significantly alleviates forgetting. Additionally, our method outperforms other methods among the three metrics. It has been discussed in Section V-D that Herding may not be an appropriate sampling strategy for Continual-AQA. Hence, we further adopt our Grouping Sampling on PODNet and AFC for better comparison. As shown in Table IX, when our Grouping Sampling is adopted, existing CL methods can perform better but still underperform our approach, further showing our approach’s effectiveness. Moreover, we calculated the parameters needed for training of different methods and revised TABLE VIII in our manuscript. As shown in Table IX, we can conclude that our method achieves significant improvement with only 1.17% increase in the number of parameters (from 1.265m to 1.279m). Note that in experiments on AQA-7, the parameter increase mainly comes from two aspects. First, the Action-Specific Graphs bring nearly 13.9k additional parameters (less than 2k parameters for each task), and the difference regressor only brings nearly 1.03k additional parameters.
TABLE XI: Comparisons of our approach with existing Continual Learning methods on the BEST dataset. “w/ Our GS” indicates our Grouping Sampling is adopted to sample the exemplars.
@@
- Robustness study on memory size. We further evaluate the robustness of our approach by adjusting the total memory size (M). As shown in Table X, our approach outperforms PODNet [26] regardless of the memory size, even if our GS is adopted.
TABLE XII: Comparisons of our approach with existing Continual Learning methods on the MTL-AQA dataset. “w/ Our GS” indicates our Grouping Sampling is adopted to sample the exemplars.
@@
Figure 7: Visualizations of the performance matrix $P$ learned by EWC [24], iCaRL [18], PODNet [26], our naive model, our GS, our FSCAR and our full model on AQA-7 dataset. The i-row-j-column element of $P$ denotes the Spearman’s rank correlation coefficient on the $i^{th}$ task after learning the $j^{th}$ task. Darker color indicates better performance. The task sequence is: Sync.10m, Skiing, Gym vault, Snowboard, Diving, Sync.3m.